40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Feb 25 15:30:29 2013, yuta, Update, Alignment, Y arm locked, both colors 
    Reply  Mon Feb 25 17:58:28 2013, rana, Update, Alignment, Y arm locked, both colors 
       Reply  Mon Feb 25 22:42:32 2013, yuta, Summary, Alignment, current IFO situation IPANGdrift.png
    Reply  Mon Feb 25 21:25:14 2013, yuta, Update, Alignment, both arms locked in IR QUAD1_1045890588.png
       Reply  Tue Feb 26 01:49:27 2013, Jenne, Update, Locking, PRMI locked 
          Reply  Tue Feb 26 02:04:51 2013, Koji, Update, Locking, PRMI locked 
             Reply  Tue Feb 26 11:55:11 2013, Manasa, Update, Locking, FPMI locked 
                Reply  Tue Feb 26 15:32:47 2013, Jenne, Update, Locking, FPMI locked 
             Reply  Wed Feb 27 01:09:53 2013, yuta, Update, Locking, PRMI sideband/carrier locked 6x
                Reply  Wed Feb 27 02:44:39 2013, Jenne, Update, Locking, DRFPMI flashes 
                   Reply  Wed Feb 27 20:10:43 2013, rana, Update, Locking, DRFPMI flashes 
Message ID: 8179     Entry time: Wed Feb 27 02:44:39 2013     In reply to: 8177     Reply to this: 8191
Author: Jenne 
Type: Update 
Category: Locking 
Subject: DRFPMI flashes 

Before Yuta left, I asked him to restore all optics to last saved values, to avoid hysteresis.

Some interesting modes appear at ETMYT and ETMXT.  The cavities aren't super well aligned right now, especially since we have been seeing this input pointing drift, but it's cool to see our first DRFPMI flashes. SRM, in particular, hasn't been aligned since before pumpdown.  If I misalign SRM, the mode content at the trans cameras improves somewhat, but it still isn't all low-order modes.

 

Here is the note that I wrote on youtube describing the video:

DRFPMI Flashing

Upper left is AS, upper right is POP
lower left is Xarm Trans (ETMXT), lower right is Yarm Trans (ETMYT)

Arms were last aligned several hours ago, and we know input pointing drifts, so at least some of the higher order modes in the arm transmissions are due to poor input pointing. All optics are "restored" to their last saved values, including SRM, which has not been aligned since before pumpdown.

TRX DC PD values are flashing to as high as 10
TRY DC PD values are flashing to as high as 7

Since uploading the video, I have seen TRY flash (once) to 45.  Yes, forty five!!!  Both arms are usually flashing to ~3ish, with an occasional medium flash (5-10), and then a few rare TRY flashes above 20.

We may have officially lost the bet of having arms locked by 9am Monday, but I think Team Grad Student / Postdoc still deserves some beer from Team Faculty / Staff.

ELOG V3.1.3-